# Adversarial Reinforcement Learning Framework for ESP Cheater Simulation

### v1

@@ A.1 Gridworld
Gridworld environment is a two-dimensional grid world based on the MiniGrid (Chevalier-Boisvert et al., 2023) framework. It contains $n\times n$ tiles, surrounded by walls that prevent the agent from leaving the environment. At the beginning of each episode, we randomly place the agent with the random facing direction. We also randomly distribute the collectible items, the walls, and the lava across empty cells. Fig. 8(a) shows the example of the initial state and the initial observation for the agents. The episode ends when the agent has collected all items or the maximum number of timesteps has elapsed.
@@
The cheater and the non-cheater receive observation dictionaries consisting of five components: Agent view, Movement history, Retrospective board, Item, and Time. Agent view provides a top-down visual representation of the environment from the agent’s perspective at the current timestep. Only a small square region in front of the agent is visible to them. Movement history is represented as a heatmap that records the agent’s movement over time; each cell stores the timestep at which the agent last visited that location, with unvisited cells set to zero. We normalize the recorded timestep values to the range $[0,1]$, where higher values correspond to more recent visits. Fig. 8(c) is an example of the movement history heatmap. Retrospective board (Gao et al., 2019) is a global memory map that records the latest observed information for each tile. Whenever a tile becomes visible to the agent, the retrospective board is updated accordingly. Non-cheater agents can only access their own partially constructed retrospective board based on their observations, whereas cheater agents have access to the full environment without any visibility restrictions. Fig. 8(b) and 8(d) illustrate examples of the retrospective board constructed by the non-cheater and the cheater agent. Item denotes the number of items collected by the agent, and Time indicates the current timestep within the episode.
@@ Appendix B Experiment Setting Details
We conducted all experiments on a machine with AMD EPYC 7742 (64-core), NVIDIA A100 (40GB) and 128GB of RAM. We use a CNN-based actor-critic architecture for the policy network and a CNN-based classifier for the detector network. Both the cheater and non-cheater policies are trained using Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017), with Generalized Advantage Estimation (GAE) (Schulman et al., 2015). We set the GAE parameter to 0.95, the clipping range to 0.2, and the value function loss coefficient to 0.5. We perform the training using 64 parallel environments, each generating rollouts of 2048 steps. PPO updates are applied over 4 epochs with a minibatch size of 1024. An entropy coefficient of 0.01 is used to encourage exploration. As for the detector, it is trained with a batch size of 8. We optimize both the policy and the detector networks using Adam optimizer (Kingma & Ba, 2014) with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a learning rate of $3\times 10^{-4}$.
@@ Appendix D Effect of Detector Design
Feature engineering (i.e., selecting important features) plays a crucial role in cheat detection (Alayed et al., 2013). Detector performance can be changed significantly depending on the selected features. Previously, we used the trajectory as a feature to implement the detector. Alternatively, we could use different features such as trajectory length and reward.

### v2

@@ A.1 Gridworld
Gridworld environment is a two-dimensional grid world based on the MiniGrid (MinigridMiniworld23) framework. It contains $n\times n$ tiles, surrounded by walls that prevent the agent from leaving the environment. At the beginning of each episode, we randomly place the agent with the random facing direction. We also randomly distribute the collectible items, the walls, and the lava across empty cells. Fig. 8(a) shows the example of the initial state and the initial observation for the agents. The episode ends when the agent has collected all items or the maximum number of timesteps has elapsed.
@@
The cheater and the non-cheater receive observation dictionaries consisting of five components: Agent view, Movement history, Retrospective board, Item, and Time. Agent view provides a top-down visual representation of the environment from the agent’s perspective at the current timestep. Only a small square region in front of the agent is visible to them. Movement history is represented as a heatmap that records the agent’s movement over time; each cell stores the timestep at which the agent last visited that location, with unvisited cells set to zero. We normalize the recorded timestep values to the range $[0,1]$, where higher values correspond to more recent visits. Fig. 8(c) is an example of the movement history heatmap. Retrospective board (gao2019skynet) is a global memory map that records the latest observed information for each tile. Whenever a tile becomes visible to the agent, the retrospective board is updated accordingly. Non-cheater agents can only access their own partially constructed retrospective board based on their observations, whereas cheater agents have access to the full environment without any visibility restrictions. Fig. 8(b) and 8(d) illustrate examples of the retrospective board constructed by the non-cheater and the cheater agent. Item denotes the number of items collected by the agent, and Time indicates the current timestep within the episode.
@@ Appendix B Experiment Setting Details
We conducted all experiments on a machine with AMD EPYC 7742 (64-core), NVIDIA A100 (40GB) and 128GB of RAM. We use a CNN-based actor-critic architecture for the policy network and a CNN-based classifier for the detector network. Both the cheater and non-cheater policies are trained using Proximal Policy Optimization (PPO) algorithm (schulman2017proximal), with Generalized Advantage Estimation (GAE) (schulman2015high). We set the GAE parameter to 0.95, the clipping range to 0.2, and the value function loss coefficient to 0.5. We perform the training using 64 parallel environments, each generating rollouts of 2048 steps. PPO updates are applied over 4 epochs with a minibatch size of 1024. An entropy coefficient of 0.01 is used to encourage exploration. As for the detector, it is trained with a batch size of 8. We optimize both the policy and the detector networks using Adam optimizer (kingma2014adam) with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a learning rate of $3\times 10^{-4}$.
@@ Appendix D Effect of Detector Design
Feature engineering (i.e., selecting important features) plays a crucial role in cheat detection (alayed2013behavioral). Detector performance can be changed significantly depending on the selected features. Previously, we used the trajectory as a feature to implement the detector. Alternatively, we could use different features such as trajectory length and reward.
@@
Appendix E Theoretical Relationship Between the Cheater Policy and the Detection Performance
In this section, we study how the choice of the cheater policy affects the detection performance.
Relationship between Cheater Policy and KL Divergence.
KL divergence between the cheater policy $\pi_{c}$ and the non-cheater policy $\pi_{n}$:
[…]
where the inequality follows from Jensen’s inequality. Let $q=\mathbb{E}_{(s,o)\sim\mathcal{S}\times\Omega}\mathbb{E}_{a_{c}\sim\pi_{c}}[\pi_{n}(a_{c}|s,o)]$ and $q^{*}=\mathbb{E}_{o\sim\Omega}[\max_{a\in\mathcal{A}}\pi_{c}(a|o)]$. Then,
$\begin{split}q=\mathbb{E}_{(s,o)\sim\mathcal{S}\times\Omega}\mathbb{E}_{a_{c}\sim\pi_{c}}[\pi_{n}(a_{c}|s,o)]&=\mathbb{E}_{(s,o)\sim\mathcal{S}\times\Omega}\sum_{a\in\mathcal{A}}\pi_{c}(a|s,o)\pi_{n}(a|s,o)\\ &=\mathbb{E}_{(s,o)\sim\mathcal{S}\times\Omega}\sum_{a\in\mathcal{A}}\pi_{c}(a|s,o)\pi_{n}(a|o)\\ &=\mathbb{E}_{o\sim\Omega}[\mathbb{E}_{s\sim\mathcal{S}|o}[\sum_{a\in\mathcal{A}}\pi_{c}(a|s,o)\pi_{n}(a|o)|o]]\\ &=\mathbb{E}_{o\sim\Omega}[\sum_{a\in\mathcal{A}}\pi_{n}(a|o)\mathbb{E}_{s\sim\mathcal{S}|o}[\pi_{c}(a|s,o)|o]]\\ &=\mathbb{E}_{o\sim\Omega}[\sum_{a\in\mathcal{A}}\pi_{n}(a|o)\sum_{s\in\mathcal{S}}p(s|o)\pi_{c}(a|s,o)]\\ &=\mathbb{E}_{o\sim\Omega}[\sum_{a\in\mathcal{A}}\pi_{n}(a|o)\pi_{c}(a|o)]\\ &\leq\mathbb{E}_{o\sim\Omega}[\max_{a\in\mathcal{A}}\pi_{c}(a|o)]=q^{*}\end{split}$ (10)
holds. By Fano’s inequality (fano1952class),
$\begin{split}1-q^{*}\geq\frac{H(\mathcal{A}_{c}|\Omega)-\log 2}{\log|\mathcal{A}_{c}|}=\frac{H(\mathcal{A}_{c}|\Omega)-\log 2}{\log|\mathcal{A}|}\end{split}$ (11)
holds. Moreover, using the strong data-processing inequality (pippenger2002reliable), $H(\mathcal{A}_{c}|\Omega)$ has a following lower bound:
$\begin{split}H(\mathcal{A}_{c}|\Omega)&=H(\mathcal{A}_{c})-I(\mathcal{A}_{c};\Omega)\\ &\geq H(\mathcal{A}_{c})-\eta I(\mathcal{A}_{c};\mathcal{S}\times\Omega)=(1-\eta)H(\mathcal{A}_{c})+\eta H(\mathcal{A}_{c}|\mathcal{S}\times\Omega),\end{split}$ (12)
where $\eta\in[0,1]$ is a contraction coefficient. $\eta$ quantifies how much of the mutual information between $\mathcal{A}_{c}$ and $\mathcal{S}\!\times\!\Omega$ is preserved when passing through $\Omega$. A smaller $\eta$ indicates stronger information loss. Combining the above results, we obtain the following lower bound of $D_{\mathrm{KL}}(\pi_{c}||\pi_{n})$:
$\begin{split}D_{\mathrm{KL}}(\pi_{c}||\pi_{n})&\geq-\log q-\mathbb{E}_{(s,o)\sim\mathcal{S}\times\Omega}H(\pi_{c}(\cdot|s,o))\\ &\geq-\log q^{*}-\mathbb{E}_{(s,o)\sim\mathcal{S}\times\Omega}H(\pi_{c}(\cdot|s,o))\\ &\geq-\log(1-\frac{H(\mathcal{A}_{c}|\Omega)-\log 2}{\log|\mathcal{A}|})-\mathbb{E}_{(s,o)\sim\mathcal{S}\times\Omega}H(\pi_{c}(\cdot|s,o))\\ &\geq-\log(1-\frac{(1-\eta)H(\mathcal{A}_{c})+\eta H(\mathcal{A}_{c}|\mathcal{S}\times\Omega)-\log 2}{\log|\mathcal{A}|})-H(\mathcal{A}_{c}|\mathcal{S}\times\Omega).\end{split}$ (13)
If $H(\mathcal{A}_{c}|\mathcal{S}\times\Omega)$ is sufficiently small (i.e., the cheater’s action is nearly determined by $(s,o)$), the lower bound becomes positive. In addition, $H(\mathcal{A}_{c}|\mathcal{S}\times\Omega)=H(\mathcal{A}_{c})-I(\mathcal{A}_{c};\mathcal{S}\times\Omega)\leq H(\mathcal{A}_{c})$ always holds. As a result, the lower bound of the KL divergence increases as $\eta$ decreases, which is likely to increase the KL divergence.
Relationship between KL Divergence and Detection Performance.
Consider a binary hypothesis testing problem between $H_{0}\!:\!\pi_{c}$ and $H_{1}\!:\!\pi_{n}$ based on $m$ i.i.d. samples. By the Chernoff–Stein’s Lemma (thomas2006elements), the type-II error $\beta_{m}$ under a fixed and bounded type-I error $\alpha_{m}$ satisfies
$\lim_{m\to\infty}-\tfrac{1}{m}\log\beta_{m}=D_{\mathrm{KL}}(\pi_{c}\|\pi_{n}).$ (14)
Hence, a larger $D_{\mathrm{KL}}(\pi_{c}\|\pi_{n})$ implies a lower asymptotic error, leading to improved detection performance.
Conclusion.
In this setting, the contraction coefficient $\eta$ reflects the degree of information preservation from $\mathcal{S}\!\times\!\Omega$ to $\Omega$ under the given cheater policy. A smaller $\eta$ indicates greater information loss, which may arise from the cheater policy that heavily exploits the unobserved components of the state. Such a decrease in $\eta$ results in a larger lower bound of $D_{\mathrm{KL}}(\pi_{c}||\pi_{n})$, and thus higher detection performance. To summarize,
[…]